OpenStudy (anonymous):

Create a unique example of dividing a polynomial by a monomial and provide the simplified form. Explain, in complete sentences, the two ways used to simplify this expression and how you would check your quotient for accuracy.

OpenStudy (anonymous):

@DebbieG

OpenStudy (anonymous):

can you help? @DebbieG

OpenStudy (anonymous):

(x³- 2x² + x – 1) : ( x- 2)= the quotient 3x² +4x+ 9 and reminder 17 Here is the method for calculation 3x² +4x+ 9 ------------------------ ** x-2 / 3x³- 2x² + x – 1 \ first 3x³ divide by x the result is 3x² **** 3x³ - 6x²*********** 3x² multiply by ( x-2) ------------------- * - ******substraction *******4x² + x **********divide again by x etc +4x *******4x² -8x ********** 4x multiply by x-2 --------------------* -*******substraction ....etc *********+9x - 1 *********+9x -18 -----------------------* - ************+17 ***** here is the rest because 17 can’t be divided by x you can do the same way for x²+x-2 / 2x⁵-0x⁴+ 3x³- 2x² + x – 1 \

OpenStudy (anonymous):

did that help

OpenStudy (anonymous):

is it correct? @issrabi

OpenStudy (anonymous):

i hope lol

OpenStudy (anonymous):

lolol

OpenStudy (anonymous):

my math can be a bit wrong maybe but i think it is bro

OpenStudy (anonymous):

idk its just all in there its kinda confusing lol sorry :/

OpenStudy (anonymous):

its ok sorry if i couldnt help

OpenStudy (anonymous):

it did probably...im just to stupid to read it and get it lol.. dnt have a math brain lol

OpenStudy (anonymous):

me neither i dont like math but im good at it so u know i have to do it

OpenStudy (anonymous):

@DebbieG

OpenStudy (debbieg):

Well, first off, @issrabi's polynomial long division example above is a bit confusing... at first I think it was wrong, but it's just that the problem he DID is different than the problem that he STATED. He stated (x³- 2x² + x – 1)/( x- 2)= the quotient 3x² +4x+ 9 and reminder 17 But he meant to say (3x³- 2x² + x – 1)/( x- 2)= the quotient 3x² +4x+ 9 and reminder 17 As first stated, of course, you get a diffferent result... but in any case, I would not do either by long division, I would use synthetic division (sooooo much easier).... ... but that's neither here not there, since that has nothing do with your problem: \(\Large \text{Create a unique example of dividing a polynomial}\) \(\Large \color{red}{ \text{ by a monomial}} \text{ and provide the simplified form.}\) x - 2 is not a monomial, it is a binomial. A monomial is a one-term polynomial, and every monomial IS a polynomial (but not every polynomial is a monomial). A polynomial is just a sum of terms that have coefficients and non-negative integer exponents (which can include a constant term). so: \(x^4+3x^2+5x-4\) is a polynomial (with 4 terms) \(3x^2+5x-4\) is a polynomial (with 3 terms, called a trinomial) \(3x^2-1\) is a polynomial (with 2 terms, called a binomial) \(5x^6\) is a monomial, which is simply a polynomial with 1 term So, I suppose, your question is asking you to do something like: \[\Large \dfrac{ 4x^4+6x^3-8x^2+4x }{ 2x }\] That is "dividing a polynomial by a monomial ".... But I'm honestly not sure what it is getting at when it says " the two ways used to simplify this expression"?? I guess you might say that one way it to factor the GCF from the numerator and cancel, and another way is to break it up into a sum of individual terms, and then reduce each term?? Eg: \[\Large \dfrac{ 4x^4+6x^3-8x^2+4x }{ 2x }= \dfrac{ 4x^4}{ 2x }+ \dfrac{ 6x^3}{ 2x }- \dfrac{ 8x^2 }{ 2x }+ \dfrac{4x }{ 2x }\] ????....... not sure.

OpenStudy (anonymous):

what you did is correct by the looks it kinda looks like the example in the book

OpenStudy (anonymous):

@DebbieG

OpenStudy (anonymous):

just confused on what i would put